OpenStudy (anonymous):

Which of the following are the x and y intercepts of the equation 2x -2y = 3? A. (2, 0), (0, 2) B. (1.5, 0), (0, -1.5) C. (2/3, 0), (0, -2/3) D. (3, 0), (0, -3)

OpenStudy (anonymous):

Put the x -value as 0 to solve for y Put the y-value as 0 to solve for x

OpenStudy (anonymous):

can you explain that a little better please

OpenStudy (anonymous):

Ok I'll help find x. You change the y into 0 to find the value of x: 2x -2y = 3 2x - 2(0) = 3 2x = 3 /2 x = 3/2 = 1.5 So your x-intercept is 1.5 Do you understand what I mean?

OpenStudy (anonymous):

(1.5, 0)

OpenStudy (anonymous):

ok thanks